# Not Just for Archiving: Provable Benefits of Reusing the Archive in Evolutionary Multi-objective Optimization

### v1

@@ Abstract
Evolutionary Algorithms (EAs) have become the most popular tool for solving widely-existed multi-objective optimization problems. In Multi-Objective EAs (MOEAs), there is increasing interest in using an archive to store non-dominated solutions generated during the search. This approach can 1) mitigate the effects of population oscillation, a common issue in many MOEAs, and 2) allow for the use of smaller, more practical population sizes. In this paper, we analytically show that the archive can even further help MOEAs through reusing its solutions during the process of new solution generation. We first prove that using a small population size alongside an archive (without incorporating archived solutions in the generation process) may fail on certain problems, as the population may remove previously discovered but promising solutions. We then prove that reusing archive solutions can overcome this limitation, resulting in at least a polynomial speedup on the expected running time. Our analysis focuses on the well-established SMS-EMOA algorithm applied to the commonly studied OneJumpZeroJump problem as well as one of its variants. We also show that reusing archive solutions can be better than using a large population size directly. Finally, we show that our theoretical findings can generally hold in practice by experiments on four well-known practical optimization problems – multi-objective 0-1 Knapsack, TSP, QAP and NK-landscape problems – with realistic settings.
@@ Introduction
Multi-objective optimization deals with problems in which multiple, often conflicting, objectives must be optimized simultaneously. Unlike single-objective optimization, there is generally no single optimal solution for a Multi-objective Optimization Problem (MOP). Instead, the goal is to find a set of solutions representing different optimal trade-offs among the objectives, called Pareto optimal solutions. The corresponding objective vectors of these solutions constitute the Pareto front. Thus, the goal of multi-objective optimization is to find the Pareto front or its good approximation. Evolutionary Algorithms (EAs), a class of randomized heuristic optimization algorithms inspired by natural evolution, have proven to be well-suited for solving MOPs due to their population-based nature. They have been widely applied across various real-world domains (Deb 2001; Zhou, Yu, and Qian 2019). Over the years, numerous well-established Multi-Objective EAs (MOEAs) have been developed, including the Non-dominated Sorting Genetic Algorithm II (NSGA-II) (Deb et al. 2002), the Multi-Objective Evolutionary Algorithm based on Decomposition (MOEA/D) (Zhang and Li 2007), and the $\mathcal{S}$-Metric Selection Evolutionary Multi-objective Optimization Algorithm (SMS-EMOA) (Beume, Naujoks, and Emmerich 2007).
@@
Original algorithm $O(\mu n^{k})$ (Bian et al. 2025) $O(\mu\cdot\max\{n^{a}/k,n^{k-a}\})$ (Theorem 5)
@@
In the area of MOEAs, there has recently been growing interest in using an archive to store non-dominated solutions generated during the search process (Li, López-Ibáñez, and Yao 2024). The main reason for this practice is as follows. In MOPs, there can be infinite many Pareto optimal solutions (Figueira et al. 2017), and it may not be practical to use a very large population aiming for accommodating all of them. However, using a small population may easily lead to the oscillation phenomenon, even when an elite preservation mechanism is employed (Knowles and Corne 2003b). This is because, during the evolutionary process, the number of generated non-dominated solutions usually exceeds the population size. As a result, population maintenance becomes necessary to remove excess non-dominated solutions (e.g., based on criteria of crowding distance). Yet, the algorithm may later accept new solutions that are non-dominated w.r.t. the current population but are actually dominated by the previously discarded solutions. Consequently, MOEAs may end up with a final population containing a large number of sub-optimal (i.e., globally dominated) solutions. This issue affects all practical MOEAs (Li, López-Ibáñez, and Yao 2024). For example, Li and Yao (2019) reports that on some problems like DTLZ7 (Deb et al. 2005), nearly half of the final population produced by NSGA-II and MOEA/D are not globally optimal (i.e., being dominated by previously discarded solutions). This phenomenon leads to the popularity of using an archive in MOEAs that store non-dominated solutions found during the search process, e.g., (Fieldsend, Everson, and Singh 2003; Krause et al. 2016; Brockhoff and Tušar 2019; Ishibuchi, Pang, and Shang 2020). Recently, theoretical studies (Bian et al. 2024) also confirm this practice, showing that using an archive can provide a speedup, particularly when working with stochastic population updates in MOEAs (Ren et al. 2025).
@@
  • Reusing the archive with a small population can be better than directly using a large population size. By comparing the results in the last two rows of Table 1, we show that for SMS-EMOA solving OneJumpZeroJump and OneJumpZeroJumpSS, reusing the archive can reduce the upper bound on the expected running time compared to using a large population size. The reason is that reusing the archive allows for a constant population size $\mu\geq 2$, which brings a speedup of $\Theta(n)$. On OneJumpZeroJumpSS, this improvement becomes evident only when $a>k/2$. Actually, Bian et al. (2024) showed that compared to using a large population size, using an archive can provide a polynomial-speedup for SMS-EMOA solving OneMinMax and LeadingOnesTrailingZeroes. In this paper, we also show that reusing the archive does not diminish this advantage.
We also conduct experiments on the artificial problems studied in this paper and four well-known practical problems: the multi-objective 0/1 knapsack (Teghem 1994), TSP (Ribeiro et al. 2002), QAP (Knowles and Corne 2003a), and NK-landscapes (Aguirre and Tanaka 2004). The results validate our theoretical findings.
It is worth pointing out that there have been many important studies on the running time of MOEAs over the past two decades. Early theoretical works (Laumanns, Thiele, and Zitzler 2004a; Laumanns, Thiele, and Zitzler 2004b; Neumann 2007; Giel and Lehre 2010; Neumann and Theile 2010; Doerr, Kodric, and Voigt 2013; Qian, Yu, and Zhou 2013; Qian, Tang, and Zhou 2016; Bian, Qian, and Tang 2018) mainly focus on simple MOEAs like (G)SEMO. Recently, researchers have begun to examine practical MOEAs. Huang et al. (2021) investigated MOEA/D. Zheng et al. (2022) conducted the first theoretical analysis of NSGA-II. Bian et al. (2025) analyzed the running time of SMS-EMOA and showed that stochastic population update can bring acceleration. Wietheger and Doerr (2023) showed that NSGA-III (Deb and Jain 2014) outperforms NSGA-II on 33OneMinMax. Ren et al. (2024a) analyzed SPEA2. Some other works on well-established MOEAs include (Zheng and Doerr 2024a; Cerf et al. 2023; Dang et al. 2023a, b; Doerr and Qu 2023b, c; Dang, Opris, and Sudholt 2024; Opris, Dang, and Sudholt 2024; Ren et al. 2024b; Wietheger and Doerr 2024; Doerr, Krejca, and Rudolph 2025; Doerr, Ivan, and Krejca 2025; Opris 2025).
@@
Multi-objective optimization aims to optimize two or more objective functions simultaneously, as shown in Definition 1. In this paper, we focus on maximization, and minimization can be defined analogously. In multi-objective optimization, objectives are typically conflicting, meaning that there is no canonical complete order in the solution space $\mathcal{X}$. To compare solutions, we use the domination relationship in Definition 2. A solution is Pareto optimal if no other solution in $\mathcal{X}$ dominates it. The set of objective vectors corresponding to all Pareto optimal solutions forms the Pareto front. The goal of multi-objective optimization is to find the Pareto front or its good approximation.
@@
  • $\bm{x}$ weakly dominates $\bm{y}$ (denoted as $\bm{x}\succeq\bm{y}$) if for any $1\leq i\leq m,f_{i}(\bm{x})\geq f_{i}(\bm{y})$;
@@
Note that the notions of “weakly dominate” and “dominate” are also called “dominate” and “strictly dominate” in some works (Cerf et al. 2023; Wietheger and Doerr 2023).
@@
SMS-EMOA (Beume, Naujoks, and Emmerich 2007) presented in Algorithm 1 is a popular MOEA, which employs non-dominated sorting and hypervolume indicator to update the population. It starts with an initial population of $\mu$ solutions (line 1). In each generation, it randomly selects a solution $\bm{x}$ from the current population (line 3) for reproduction. Afterwards, bit-wise mutation flips each bit of $\bm{x}$ with probability $1/n$ to produce an offspring $\bm{x}^{\prime}$ (line 4). Then, the union of the current population and the offspring solution is partitioned into non-dominated sets $R_{1},\ldots,R_{v}$ (line 5), where $R_{1}$ contains all non-dominated solutions in $P\cup\{\bm{x}^{\prime}\}$, and $R_{i}$ ($i\geq 2$) contains all non-dominated solutions in $P\cup\{\bm{x}^{\prime}\}\setminus\cup_{j=1}^{i-1}R_{j}$. A solution $\bm{z}\in R_{v}$ is then removed (lines 6–7) by minimizing $\Delta_{\bm{r}}(\bm{x},R_{v}):=HV_{\bm{r}}(R_{v})-HV_{\bm{r}}(R_{v}\setminus\{\bm{x}\})$, where $HV_{\bm{r}}(X)$ denotes the hypervolume of the solution set $X$ with respect to a reference point $\bm{r}\in\mathbb{R}^{m}$ ($\forall i,r_{i}\leq\min_{\bm{x}\in\mathcal{X}}f_{i}(\bm{x})$), i.e., the volume between the reference point and the objective vectors of the solution set. A larger hypervolume indicates better approximation of the Pareto front in terms of convergence and diversity. For bi-objective problems, as defined in the original SMS-EMOA (Beume, Naujoks, and Emmerich 2007), the algorithm omits the reference point and directly preserves the two boundary points, i.e., the two solutions with the maximum $f_{1}$ and $f_{2}$ values, respectively, allowing the hypervolume contribution to be calculated accordingly.
@@
Next, we introduce benchmark problems OneJumpZeroJump and its variant OneJumpZeroJumpSS, studied in this paper. The OneJumpZeroJump problem as presented in Definition 3 is constructed based on the Jump problem (Doerr and Neumann 2020), and has been widely used in MOEAs’ theoretical analyses (Doerr and Zheng 2021; Doerr and Qu 2023a; Lu, Bian, and Qian 2024; Ren et al. 2024b). Its first objective is the same as the Jump problem, which aims to maximize the number of 1-bits of a solution, except for a valley around $1^{n}$ (the solution with all 1-bits) where the number of 1-bits should be minimized. The second objective is isomorphic to the first, with the roles of 1-bits and 0-bits reversed. The Pareto front of the OneJumpZeroJump problem is $\{(c,n+2k-c)\mid c\in[2k..n]\cup\{k,n+k\}\}$, whose size is n2k+3n-2k+3, and the Pareto optimal solution corresponding to (c,n+2kc)(c,n+2k-c) is any solution with (ck)(c-k) 1-bits. We use $S_{I}^{*}=\{\bm{x}\mid|\bm{x}|_{1}\in[k..n-k]\}$ and $F_{I}^{*}=\{(c,n+2k-c)\mid c\in[2k..n]\}$ to denote the internal part of Pareto set and Pareto front, respectively.
Definition 3 (Doerr and Zheng (2021)).
@@
$f_{1}(\bm{x})=\begin{cases}2k+1/n,&\text{if }|\bm{x}|_{1}=k-a,\\ n-1/n,&\text{else if }|\bm{x}|_{1}=n-(k-a),\\ k+|\bm{x}|_{1},&\text{else if }|\bm{x}|_{1}\leq n-k\text{ or }\bm{x}=1^{n},\\ n-|\bm{x}|_{1},&\text{else},\end{cases}$
$f_{2}(\bm{x})=f_{1}(\bar{\bm{x}}),$
@@
Lemma 1 (Robbins (1955)).
@@
Conditioned on the occurrence of $A_{1}$, let $A_{2}$ denote the event that the central point on the Pareto front (i.e., solutions with $n/2$ 1-bits for even $n$ and (n1)/2(n-1)/2 1-bits for odd $n$) can only be generated from $0^{n}$ or $1^{n}$. Then, we show that $A_{2}$ occurs with probability at least $e^{-8e(e-1)}\cdot e^{-1/\Omega(n)}$. We note that in the original SMS-EMOA (Beume, Naujoks, and Emmerich 2007), when solving bi-objective problems, the algorithm directly preserves the two boundary points, i.e., the two solutions with the maximum $f_{1}$ and $f_{2}$ values respectively. Thus, the maximum $f_{1}$ and $f_{2}$ value among the Pareto optimal solutions in $P\cup\{\bm{x}^{\prime}\}$ will not decrease, where $\bm{x}^{\prime}$ is the offspring solution generated in each generation. Assume that $n$ is even and the solution with the maximum $f_{2}$ value in the population is $\bm{x}^{*}$. According to the definition of OneJumpZeroJump in Def 3, $\bm{x}^{*}$ has the minimum number of 1-bits. Suppose that in some generation, $\bm{x}^{*}$ such that $|\bm{x}^{*}|_{1}=i$, is selected as the parent solution. By flipping a single 1-bits with probability $(i/n)\cdot(1-1/n)^{n-1}\geq i/(en)$, the maximum $f_{2}$ value of population increases. Or by flipping n/2in/2-i 0-bits with probability $(1-1/n)^{n/2-i}\cdot\binom{n-i}{n/2-i}/n^{n/2-i}\leq\binom{n-i}{n/2-i}/(n^{n/2-i})$, a solution with $n/2$ 1-bits is found. Thus, if choosing $\bm{x}^{*}$ as the parent, the conditional probability of increasing $f_{2}$ value rather than finding a solution with $n/2$ 1-bits is at least
@@
For the first phase, we prove that the expected number of generations for finding one point on $F_{I}^{*}$ is $O(1)$. During population initialization, by the Chernoff bound and $n-2k=\Omega(n)$, all solutions in the initial population has at most k1k-1 1-bits or at least nk+1n-k+1 1-bits with probability at most $\exp(-\Omega(n))^{\mu}$. Without loss of generality, we assume that one solution $\bm{y}$ has at most k1k-1 1-bits. Then similar to the analysis in (Bian et al. 2025), selecting $\bm{y}$ and flipping $k-|\bm{y}|_{1}$ 0-bits can generate a point on $F_{I}^{*}$, whose probability is at least
@@
For the second phase, we show that the expected number of generations for finding $1^{n}$ is $O(\mu n^{k})$, and by symmetry, the same bound holds for finding $0^{n}$. We note that in the original SMS-EMOA (Beume, Naujoks, and Emmerich 2007), when solving bi-objective problems, the algorithm omits the reference point and directly preserves the two boundary points, i.e., the two solutions with the maximum $f_{1}$ and $f_{2}$ values respectively (with ties broken randomly in the case of identical objective vectors), and computes the hypervolume contribution of the remaining solutions based on these two points. Thus, the maximum $f_{1}$ value among the Pareto optimal solutions in $P\cup\{\bm{x}^{\prime}\}$ will not decrease, where $\bm{x}^{\prime}$ is the offspring solution generated in each generation. We assume that $1^{n}$ has not been found; otherwise, the second phase finishes. Now, we consider the increase of the maximum $f_{1}$ value among the Pareto optimal solutions found. In each generation, SMS-EMOA chooses the population $P$ as the parent population with probability $1/2$, the solution $\bm{x}$ with the maximum $f_{1}$ value is selected with probability $1/\mu$, and generates a solution with more 1-bits that only one of the 0-bits is flipped by bit-wise mutation with probability $((n-|\bm{x}|_{1})/n)\cdot(1-1/n)^{n-1}$. This implies that the probability of generating a solution with more than $|\bm{x}|_{1}$ 1-bits in one generation is at least
@@
In the last section, we theoretically show that reusing the archive is beneficial compared to using an archive without reuse. Now, we show that reusing the archive with a small population can be better than directly using a large population. Specifically, we further show that when reusing the archive, the upper bound on the expected running time for SMS-EMOA solving OneJumpZeroJumpSS, OneJumpZeroJump, OneMinMax, and LeadingOnesTrailingZeroes is lower than that of using a large population size. We prove in Theorem 5 that the expected running time for SMS-EMOA solving OneJumpZeroJumpSS is $O(\mu\cdot\max\{n^{a}/k,n^{k-a}\})$ if using a population size $\mu\geq n-2k+5$. The proof idea is to divide the optimization procedure into three phases, where the first phase aims at finding the whole internal Pareto front $F_{I}^{*}$, the second phase aims at finding the two points on $G$, and the third phase aims at finding $1^{n}$ and $0^{n}$. In Theorem 6, Bian et al. (2025) showed that the expected running time for SMS-EMOA solving OneJumpZeroJump is $O(\mu n^{k})$ if using a population size $\mu\geq n-2k+3$.
@@
Theorem 6 (Bian et al. (2025)).
@@
Bian et al. (2024) showed that for SMS-EMOA solving OneMinMax and LeadingOnesTrailingZeroes, if using a small population and an archive without reuse, the upper bound on the expected running time can be reduced by a factor of $\Theta(n)$, compared to using a large population size $\mu\geq n+1$ (i.e., the size of Pareto front). Next, we will show that archive reuse does not diminish this advantage. Note that Bian et al. (2024) considered one-point crossover, so the archive reuse method is adapted to select each parent solution for recombination with probability $1/2$ from the population and $1/2$ from the archive. Then, we only need to consider the case where both parents are selected from the population (whose probability is $1/4$), and the remaining analysis is the same as that in (Bian et al. 2024). The definition of OneMinMax and LeadingOnesTrailingZeroes, and full proofs are given in the supplementary.
Theorem 7 (Zheng et al. (2024b)).
@@
The expected running time for SMS-EMOA (without using an archive) solving OneMinMax and LeadingOnesTrailingZeroes has been shown to be $O(\mu n\log n)$ and $O(\mu n^{2})$ in Theorems 7 and 9, respectively, where the population size $\mu\geq n+1$ (Zheng and Doerr 2024b). Thus, our results in Theorems 8 and 10 show that if a constant population size is used for SMS-EMOA reusing the archive, the expected running time can be reduced by a factor of $\Theta(n)$.

### v2

@@ Abstract
Evolutionary Algorithms (EAs) have become the most popular tool for solving widely-existed multi-objective optimization problems. In Multi-Objective EAs (MOEAs), there is increasing interest in using an archive to store non-dominated solutions generated during the search. This approach can 1) mitigate the effects of population oscillation, a common issue in many MOEAs, and 2) allow for the use of smaller, more practical population sizes. In this paper, we analytically show that the archive can even further help MOEAs through reusing its solutions during the process of new solution generation. We first prove that using a small population size alongside an archive (without incorporating archived solutions in the generation process) may fail on certain problems, as the population may remove previously discovered but promising solutions. We then prove that reusing archive solutions can overcome this limitation, resulting in at least a polynomial speedup on the expected running time. Our analysis focuses on the well-established SMS-EMOA algorithm applied to the commonly studied OneJumpZeroJump problem as well as one of its variants. We also show that reusing archive solutions can be better than using a large population size directly. Finally, we show that our theoretical findings can generally hold in practice by experiments on well-known practical optimization problems – multi-objective 0-1 Knapsack, TSP, QAP and NK-landscape problems – with realistic settings.
@@ Introduction
Multi-objective optimization deals with problems in which multiple, often conflicting, objectives must be optimized simultaneously. Unlike single-objective optimization, there is generally no single optimal solution for a Multi-objective Optimization Problem (MOP). Instead, the goal is to find a set of solutions representing different optimal trade-offs among the objectives, called Pareto optimal solutions. The corresponding objective vectors of these solutions constitute the Pareto front. Thus, the goal of multi-objective optimization is to find the Pareto front or its good approximation. Evolutionary Algorithms (EAs), a class of randomized heuristic optimization algorithms inspired by natural evolution, have proven to be well-suited for solving MOPs due to their population-based nature. They have been widely applied across various real-world domains (deb2001book; qian19el). Over the years, numerous well-established Multi-Objective EAs (MOEAs) have been developed, including the Non-dominated Sorting Genetic Algorithm II (NSGA-II) (deb-tec02-nsgaii), the Multi-Objective Evolutionary Algorithm based on Decomposition (MOEA/D) (zhang2007moea), and the $\mathcal{S}$-Metric Selection Evolutionary Multi-objective Optimization Algorithm (SMS-EMOA) (beume2007sms).
@@
Original algorithm $O(\mu n^{k})$ (bian23stochastic) $O(\mu\cdot\max\{n^{a}/k,n^{k-a}\})$ (Theorem 5)
@@
In the area of MOEAs, there has recently been growing interest in using an archive to store non-dominated solutions generated during the search process (li2023multi). The main reason for this practice is as follows. In MOPs, there can be infinite many Pareto optimal solutions (figueira2017easy), and it may not be practical to use a very large population aiming for accommodating all of them. However, using a small population may easily lead to the oscillation phenomenon, even when an elite preservation mechanism is employed (knowles2003properties). This is because, during the evolutionary process, the number of generated non-dominated solutions usually exceeds the population size. As a result, population maintenance becomes necessary to remove excess non-dominated solutions (e.g., based on criteria of crowding distance). Yet, the algorithm may later accept new solutions that are non-dominated w.r.t. the current population but are actually dominated by the previously discarded solutions. Consequently, MOEAs may end up with a final population containing a large number of sub-optimal (i.e., globally dominated) solutions. This issue affects all practical MOEAs (li2023multi). For example, Li and Yao (li2019empirical) reports that on some problems like DTLZ7 (deb2005scalable), nearly half of the final population produced by NSGA-II and MOEA/D are not globally optimal (i.e., being dominated by previously discarded solutions). This phenomenon leads to the popularity of using an archive in MOEAs that store non-dominated solutions found during the search process, e.g., (fieldsend2003using; krause2016unbounded; brockhoff2019benchmarking; ishibuchi2020new). Recently, theoretical studies (bian2024archive) also confirm this practice, showing that using an archive can provide a speedup, particularly when working with stochastic population updates in MOEAs (ren2025archive).
@@
  • Reusing the archive with a small population can be better than directly using a large population size. By comparing the results in the last two rows of Table 1, we show that for SMS-EMOA solving OneJumpZeroJump and OneJumpZeroJumpSS, reusing the archive can reduce the upper bound on the expected running time compared to using a large population size. The reason is that reusing the archive allows for a constant population size $\mu\geq 2$, which brings a speedup of $\Theta(n)$. On OneJumpZeroJumpSS, this improvement becomes evident only when $a>k/2$. Actually, Bian et al. (bian2024archive) showed that compared to using a large population size, using an archive can provide a polynomial-speedup for SMS-EMOA solving OneMinMax and LeadingOnesTrailingZeroes. In this paper, we also show that reusing the archive does not diminish this advantage.
We also conduct experiments on the artificial problems studied in this paper and four well-known practical problems: the multi-objective 0/1 knapsack (teghem_multi-objective_1994), TSP (ribeiro2002study), QAP (knowles2003instance), and NK-landscapes (Aguirre2004). The results validate our theoretical findings.
It is worth pointing out that there have been many important studies on the running time of MOEAs over the past two decades. Early theoretical works (laumanns-nc04-knapsack; LaumannsTEC04; Neumann07; Giel10; Neumann10; doerr2013lower; Qian13; qian-ppsn16-hyper; bian2018tools) mainly focus on simple MOEAs like (G)SEMO. Recently, researchers have begun to examine practical MOEAs. Huang et al. (huang2021runtime) investigated MOEA/D. Zheng et al. (zheng2021first) conducted the first theoretical analysis of NSGA-II. Bian et al. (bian23stochastic) analyzed the running time of SMS-EMOA and showed that stochastic population update can bring acceleration. Wietheger and Doerr (wietheger23nsgaiii) showed that NSGA-III (deb2014nsgaiii) outperforms NSGA-II on 33OneMinMax. Ren et al. (Ren2024spea2) analyzed SPEA2. Some other works on well-established MOEAs include (bian2022better; zheng2023manyobj; cerf2023first; dang2023analysing; dang2023crossover; doerr2023lower; doerr2023crossover; dang2024level; Opris2024nsgaiii; ren2024multimodel; wietheger2024near; doerr2025runtime; doerr2025speeding; opris2025first). There are also a series of theoretical studies showing that MOEAs can achieve good approximation guarantees for single-objective submodular optimization with constraints (qian2015subset; Friedrich2015sub; qian2017subset; qian2019sub)
@@
Multi-objective optimization aims to optimize two or more objective functions simultaneously, as shown in Definition 1. The objectives are typically conflicting, meaning that there is no canonical complete order in the solution space $\mathcal{X}$. To compare solutions, we use the domination relationship in Definition 2. A solution is Pareto optimal if no other solution in $\mathcal{X}$ dominates it. The set of objective vectors corresponding to all Pareto optimal solutions forms the Pareto front. The goal of multi-objective optimization is to find the Pareto front or its good approximation.
@@
  • $\bm{x}$ weakly dominates $\bm{y}$ (denoted as $\bm{x}\succeq\bm{y}$) if for all $1\leq i\leq m,f_{i}(\bm{x})\geq f_{i}(\bm{y})$;
@@
SMS-EMOA (beume2007sms) presented in Algorithm 1 is a popular MOEA, which employs non-dominated sorting and hypervolume indicator to update the population. It starts with an initial population of $\mu$ solutions (line 1). In each generation, it randomly selects a solution $\bm{x}$ from the current population (line 3) for reproduction. Afterwards, bit-wise mutation flips each bit of $\bm{x}$ with probability $1/n$ to produce an offspring $\bm{x}^{\prime}$ (line 4). Then, the union of the current population and the offspring solution is partitioned into non-dominated sets $R_{1},\ldots,R_{v}$ (line 5), where $R_{1}$ contains all non-dominated solutions in $P\cup\{\bm{x}^{\prime}\}$, and $R_{i}$ ($i\geq 2$) contains all non-dominated solutions in $P\cup\{\bm{x}^{\prime}\}\setminus\cup_{j=1}^{i-1}R_{j}$. A solution $\bm{z}\in R_{v}$ is then removed (lines 6–7) by minimizing $\Delta_{\bm{r}}(\bm{x},R_{v}):=HV_{\bm{r}}(R_{v})-HV_{\bm{r}}(R_{v}\setminus\{\bm{x}\})$, where $HV_{\bm{r}}(X)$ denotes the hypervolume of the solution set $X$ with respect to a reference point $\bm{r}\in\mathbb{R}^{m}$ ($\forall i,r_{i}\leq\min_{\bm{x}\in\mathcal{X}}f_{i}(\bm{x})$), i.e., the volume between the reference point and the objective vectors of the solution set. A larger hypervolume indicates better approximation of the Pareto front in terms of convergence and diversity. For bi-objective problems, as defined in the original SMS-EMOA (beume2007sms), the algorithm omits the reference point and directly preserves the two boundary points, i.e., the two solutions with the maximum $f_{1}$ and $f_{2}$ values, respectively, allowing the hypervolume contribution to be calculated accordingly.
@@
Next, we introduce benchmark problems OneJumpZeroJump and its variant OneJumpZeroJumpSS, studied in this paper. The OneJumpZeroJump problem as presented in Definition 3 is constructed based on the Jump problem (doerr2020theory), and has been widely used in MOEAs’ theoretical analyses (doerr2021ojzj; doerr2023nsgaojzj; lu2024imoea; ren2024multimodel). Its first objective is the same as the Jump problem, which aims to maximize the number of 1-bits of a solution, except for a valley around $1^{n}$ (the solution with all 1-bits) where the number of 1-bits should be minimized. The second objective is isomorphic to the first, with the roles of 1-bits and 0-bits reversed. The Pareto front of the OneJumpZeroJump problem is $\{(c,n+2k-c)\mid c\in[2k..n]\cup\{k,n+k\}\}$, whose size is n2k+3n-2k+3, and the Pareto optimal solution corresponding to (c,n+2kc)(c,n+2k-c) is any solution with (ck)(c-k) 1-bits. We use $S_{I}^{*}=\{\bm{x}\mid|\bm{x}|_{1}\in[k..n-k]\}$ and $F_{I}^{*}=\{(c,n+2k-c)\mid c\in[2k..n]\}$ to denote the internal part of Pareto set and Pareto front, respectively.
Definition 3 (Doerr and Zheng (doerr2021ojzj)).
@@
$f_{1}(\bm{x})=\begin{cases}2k+1/n,&\text{if }|\bm{x}|_{1}=k-a,\\ n-1/n,&\text{else if }|\bm{x}|_{1}=n-(k-a),\\ k+|\bm{x}|_{1},&\text{else if }|\bm{x}|_{1}\leq n-k\text{ or }\bm{x}=1^{n},\\ n-|\bm{x}|_{1},&\text{else},\end{cases}$ (1)
$f_{2}(\bm{x})=f_{1}(\bar{\bm{x}}),$ (2)
@@
Lemma 1 (Robbins (robbins1955remark)).
@@
Conditioned on the occurrence of $A_{1}$, let $A_{2}$ denote the event that the central point on the Pareto front (i.e., solutions with $n/2$ 1-bits for even $n$ and (n1)/2(n-1)/2 1-bits for odd $n$) can only be generated from $0^{n}$ or $1^{n}$. Then, we show that $A_{2}$ occurs with probability at least $e^{-8e(e-1)}\cdot e^{-1/\Omega(n)}$. We note that in the original SMS-EMOA (beume2007sms), when solving bi-objective problems, the algorithm directly preserves the two boundary points, i.e., the two solutions with the maximum $f_{1}$ and $f_{2}$ values respectively. Thus, the maximum $f_{1}$ and $f_{2}$ value among the Pareto optimal solutions in $P\cup\{\bm{x}^{\prime}\}$ will not decrease, where $\bm{x}^{\prime}$ is the offspring solution generated in each generation. Assume that $n$ is even and the solution with the maximum $f_{2}$ value in the population is $\bm{x}^{*}$. According to the definition of OneJumpZeroJump in Def 3, $\bm{x}^{*}$ has the minimum number of 1-bits. Suppose that in some generation, $\bm{x}^{*}$ such that $|\bm{x}^{*}|_{1}=i$, is selected as the parent solution. By flipping a single 1-bits with probability $(i/n)\cdot(1-1/n)^{n-1}\geq i/(en)$, the maximum $f_{2}$ value of population increases. Or by flipping n/2in/2-i 0-bits with probability $(1-1/n)^{n/2-i}\cdot\binom{n-i}{n/2-i}/n^{n/2-i}\leq\binom{n-i}{n/2-i}/(n^{n/2-i})$, a solution with $n/2$ 1-bits is found. Thus, if choosing $\bm{x}^{*}$ as the parent, the conditional probability of increasing $f_{2}$ value rather than finding a solution with $n/2$ 1-bits is at least
@@
For the first phase, we prove that the expected number of generations for finding one point on $F_{I}^{*}$ is $O(1)$. During population initialization, by the Chernoff bound and $n-2k=\Omega(n)$, all solutions in the initial population has at most k1k-1 1-bits or at least nk+1n-k+1 1-bits with probability at most $\exp(-\Omega(n))^{\mu}$. Without loss of generality, we assume that one solution $\bm{y}$ has at most k1k-1 1-bits. Then similar to the analysis in (bian23stochastic), selecting $\bm{y}$ and flipping $k-|\bm{y}|_{1}$ 0-bits can generate a point on $F_{I}^{*}$, whose probability is at least
@@
For the second phase, we show that the expected number of generations for finding $1^{n}$ is $O(\mu n^{k})$, and by symmetry, the same bound holds for finding $0^{n}$. We note that in the original SMS-EMOA (beume2007sms), when solving bi-objective problems, the algorithm omits the reference point and directly preserves the two boundary points, i.e., the two solutions with the maximum $f_{1}$ and $f_{2}$ values respectively (with ties broken randomly in the case of identical objective vectors), and computes the hypervolume contribution of the remaining solutions based on these two points. Thus, the maximum $f_{1}$ value among the Pareto optimal solutions in $P\cup\{\bm{x}^{\prime}\}$ will not decrease, where $\bm{x}^{\prime}$ is the offspring solution generated in each generation. We assume that $1^{n}$ has not been found; otherwise, the second phase finishes. Now, we consider the increase of the maximum $f_{1}$ value among the Pareto optimal solutions found. In each generation, SMS-EMOA chooses the population $P$ as the parent population with probability $1/2$, the solution $\bm{x}$ with the maximum $f_{1}$ value is selected with probability $1/\mu$, and generates a solution with more 1-bits that only one of the 0-bits is flipped by bit-wise mutation with probability $((n-|\bm{x}|_{1})/n)\cdot(1-1/n)^{n-1}$. This implies that the probability of generating a solution with more than $|\bm{x}|_{1}$ 1-bits in one generation is at least
@@
In the last section, we theoretically show that reusing the archive is beneficial compared to using an archive without reuse. Now, we show that reusing the archive with a small population can be better than directly using a large population. Specifically, we further show that when reusing the archive, the upper bound on the expected running time for SMS-EMOA solving OneJumpZeroJumpSS, OneJumpZeroJump, OneMinMax, and LeadingOnesTrailingZeroes is lower than that of using a large population size. We prove in Theorem 5 that the expected running time for SMS-EMOA solving OneJumpZeroJumpSS is $O(\mu\cdot\max\{n^{a}/k,n^{k-a}\})$ if using a population size $\mu\geq n-2k+5$. The proof idea is to divide the optimization procedure into three phases, where the first phase aims at finding the whole internal Pareto front $F_{I}^{*}$, the second phase aims at finding the two points on $G$, and the third phase aims at finding $1^{n}$ and $0^{n}$. In Theorem 6, Bian et al. (bian23stochastic) showed that the expected running time for SMS-EMOA solving OneJumpZeroJump is $O(\mu n^{k})$ if using a population size $\mu\geq n-2k+3$.
@@
Theorem 6 (Bian et al. (bian23stochastic)).
@@
Bian et al. (bian2024archive) showed that for SMS-EMOA solving OneMinMax and LeadingOnesTrailingZeroes, if using a small population and an archive without reuse, the upper bound on the expected running time can be reduced by a factor of $\Theta(n)$, compared to using a large population size $\mu\geq n+1$ (i.e., the size of Pareto front). Next, we will show that archive reuse does not diminish this advantage. Note that Bian et al. (bian2024archive) considered one-point crossover, so the archive reuse method is adapted to select each parent solution for recombination with probability $1/2$ from the population and $1/2$ from the archive. Then, we only need to consider the case where both parents are selected from the population (whose probability is $1/4$), and the remaining analysis is the same as that in (bian2024archive). The definition of OneMinMax and LeadingOnesTrailingZeroes, and full proofs are given in the supplementary.
Theorem 7 (Zheng et al. (zheng2024sms)).
@@
The expected running time for SMS-EMOA (without using an archive) solving OneMinMax and LeadingOnesTrailingZeroes has been shown to be $O(\mu n\log n)$ and $O(\mu n^{2})$ in Theorems 7 and 9, respectively, where the population size $\mu\geq n+1$ (zheng2024sms). Thus, our results in Theorems 8 and 10 show that if a constant population size is used for SMS-EMOA reusing the archive, the expected running time can be reduced by a factor of $\Theta(n)$.
Refer to caption Refer to caption Refer to caption
(a) KP 100D (b) KP 200D (c) KP 500D
Refer to caption Refer to caption Refer to caption
(d) NK 100D (e) NK 200D (f) NK 500D
Refer to caption Refer to caption Refer to caption
(g) TSP 100D (h) TSP 200D (i) TSP 500D
Refer to caption Refer to caption Refer to caption
(j) QAP 100D (k) QAP 200D (l) QAP 500D
Figure 2: Non-dominated solutions obtained by SMS-EMOA-A (archive only to store non-dominated solutions), SMS-EMOA-AR (archive with reuse), and SMS-EMOA-L (a large population size) on the multi-objective 0/1 Knapsack (KP), NK-Landscape ,(NK), TSP and QAP problems, with 100 (top panel), 200 (middle panel), and 500 (bottom panel) decision variables respectively. The Knapsack and NK-Landscape are maximization problems, and the TSP and QAP are minimization problems.